OpenStudy (anonymous):

Solve each equation. Show all your work. Round your answers to four decimal places. 17.) log7 (x +1) + log7(x – 5) = 1

OpenStudy (mathstudent55):

This is a rule of logs: \( \log a + \log b = \log (ab) \)

OpenStudy (mertsj):

\[\log_{7}(x-1)+\log_{7}(x+5)=1 \] \[\log_{7}(x-1)(x+5)=1 \]

OpenStudy (mertsj):

Now put that equation into its exponential form.

OpenStudy (mathstudent55):

That means change the form of the equation form the left to the right as shown below. \( log_b x = y \iff b^y = x\)

OpenStudy (anonymous):

so x^2 - 5 = 1 x^2 = -4?

OpenStudy (mertsj):

What is the base?

OpenStudy (anonymous):

log7?

OpenStudy (mertsj):

The base is 7. What is the exponent?

OpenStudy (anonymous):

how would i find that?

OpenStudy (mertsj):

The log is the exponent. You equation says: log ( )=1 That means the exponent is 1

OpenStudy (mertsj):

So you know the base and you know the exponent

OpenStudy (mertsj):

Write that. It is equal to the expression in the parentheses

OpenStudy (anonymous):

4?

OpenStudy (mertsj):

What is the base?

OpenStudy (mathstudent55):

Here's an example of how to convert from log form to exponential form.